OpenStudy (jaiv):

graph -3x^2+12y^2=84. What are the domain and range

OpenStudy (zpupster):

if you divide all terms by 84 : \[\frac{ y^2 }{ 7 } - \frac{ x^2 }{ 28 } =1\] this is the standard form of a parabola Hyperbola opening Up and down C(0,0) along x =0